# HumanCM: One Step Human Motion Prediction

### v1

@@ Abstract
We present HumanCM, a one-step human motion prediction framework built upon consistency models. Instead of relying on multi-step denoising as in diffusion-based methods, HumanCM performs efficient single-step generation by learning a self-consistent mapping between noisy and clean motion states. The framework adopts a Transformer-based spatiotemporal architecture with temporal embeddings to model long-range dependencies and preserve motion coherence. Experiments on Human3.6M and HumanEva-I demonstrate that HumanCM achieves comparable or superior accuracy to state-of-the-art diffusion models while reducing inference steps by up to two orders of magnitude.
@@ III-C HumanCM
where $w^{*}$ is a fixed guidance scale and $\lambda$ balances the two objectives. The first term enforces temporal self-consistency, while the second encourages the model to maintain semantic fidelity and motion realism in the latent space. The overall training procedure is illustrated in Figure 2, while the detailed training and inference steps are provided in Algorithm 1 and Algorithm 2, respectively.
@@ IV-B Implementation details
We train HumanCM by applying consistency distillation in a pre-trained DDIM model from Transfusion[33] with the same model structure. Experiments are conducted with PyTorch and one NVIDIA A100-SXM4-40GB GPU. Training is performed with the Adam optimizer, an initial learning rate of 0.0003 decayed by 0.9 every 75 epochs, for 5000 epochs both on Human3.6M and HumanEva-I. The training guidance scale range $w_{\min},w_{\max}]=[0,1]$ and the infer guidance scale is $w^{*}=1/125$. The scale $\lambda$ for reconstruction constraint is set as $1/15$.

### v2

@@ Abstract
We present HumanCM, a one-step human motion prediction framework built upon consistency models. Instead of relying on multi-step denoising as in diffusion-based methods, HumanCM performs efficient single-step generation by learning a self-consistent mapping between noisy and clean motion states in a latent space. By operating in this compact representation, HumanCM captures long-range temporal dependencies and preserves motion coherence. Experiments on Human3.6M and HumanEva-I demonstrate that HumanCM achieves comparable or superior accuracy to state-of-the-art diffusion models while reducing inference steps by up to two orders of magnitude.
@@ III-C HumanCM
where $w^{*}$ is a fixed guidance scale and $\lambda$ balances the two objectives. The first term enforces temporal self-consistency, while the second encourages the model to maintain semantic fidelity and motion realism in the latent space.
The overall training procedure is illustrated in Figure 2, while the detailed training and inference steps are provided in Algorithm 1 and Algorithm 2, respectively. As shown in Figure 3, HumanCM eliminates the multi-step denoising required by diffusion-based models (blue box) and performs direct one-step generation (green box), offering significant acceleration without compromising motion quality. Note that the self-attention operations are performed over latent temporal tokens after frequency transformation, rather than directly on raw joint coordinates; thus, visualizing attention maps on the spatial domain is not physically meaningful.
Figure 3: Timeline comparison of motion generation processes. The blue box illustrates the traditional diffusion-based inference, where the model progressively denoises from $y(T)$ to $y(0)$ through $T$ iterative steps. In contrast, the green box highlights our HumanCM one-step inference, which directly maps the noisy input $y(T)$ to the clean motion state $y(0)$ in a single forward pass, achieving over two orders of magnitude faster generation.
@@ IV-B Implementation details
We train HumanCM by applying consistency distillation in a pre-trained DDIM model from Transfusion[33] with the same model structure. Experiments are conducted with PyTorch and one NVIDIA A100-SXM4-40GB GPU. Training is performed with the Adam optimizer, an initial learning rate of 0.0003 decayed by 0.9 every 75 epochs, for 5000 epochs both on Human3.6M and HumanEva-I. The training guidance scale range $w_{\min},w_{\max}]=[0,1]$ and the infer guidance scale is $w^{*}=1/125$. The scale $\lambda$ for reconstruction constraint is set as $1/15$. To capture long-range temporal dependencies, HumanCM employs temporal embeddings with a fixed dimension of $d_{t}$, which is chosen to match the model’s latent dimension (e.g., 256), ensuring compatibility with the Transformer’s multi-head attention layers.